Subject: APO 3.5m users committee minutes, Nov 2012

From: Suzanne Hawley

Submitted: Wed, 28 Nov 2012 16:02:17 -0800 (PST)

Message number: 1297 (previous: 1296, next: 1298 up: Index)

 		APO 3.5-m Users Committee Phonecon, 11/05/12

Attending:  Suzanne Hawley, John Bally, Scott Anderson, Jon Holtzman, Rene 
Walterbos, Michael Strauss, Mary Beth Kaiser, Bill Ketzeback, and Bruce 

Absent:  Mark Klaene, Al Harper, Remy Indebetouw


User feedback and comments from institutional representatives:

o Chicago (Al Harper) - No report.

o Washington (Scott Anderson) - Scott had nothing new to report.

o New Mexico State (Jon Holtzman) - Jon had nothing new to report.

o Princeton (Michael Strauss) - Michael had nothing new to report.

o Colorado (John Bally) - John noted that they had a very good run last night.

o Virginia (Remy Indebetouw) - No report.

o Johns Hopkins (Mary Beth Kaiser) - Mary Beth had nothing new to report.


Discussion of telescope/instruments report:

Bill noted that we've enjoyed great observing weather recently.  The 
telescope has been solid, and the instruments have been generally OK. 
The replacement window for SPIcam is on order, and the noise from the ion 
pump appears to be a grounding issue that is being worked.  Flats taken 
with the echelle occasionally show differing levels, which is due to a 
malfunctioning flip mirror, which is proving challenging to diagnose and 
fix.  Mary Beth asked when this is likely to be resolved.  Bill said that 
several tries to fix it were so far unsuccessful, but some new parts are 
on order and he's hopeful -- otherwise we may be compelled to re-engineer 
the electronics and software for the flip mirror.  This problem has been 
with us for quite some time, and has been gradually getting worse.

Mark Klaene's detailed report for the reporting period follows:


 	      3.5-m Telescope and Instruments Highlights
 	                 09/29/12 through 10/30/12

0) Overview

Productive period with some excellent seeing at times and minimal lost time.

1) Telescope

The telescope performance has been nominal this period.

2) Instruments --

DIS has been operating nominally.  Both red and blue cameras have been 
functioning normally with good vacuum and temperature.

Agile had no issues this period.

SPIcam's vacuum leak was traced to a chip in the front window.  A 
replacement window with AR coatings is on-order.  We expect its delivery 
in the next 2 weeks.  Until then, the instrument is off-line.  An 
electrical noise issue was traced to the new ion pump.  An electrical 
insulating disc (identical to what is installed in the DIS cameras) needs 
to be installed.  The UW machine shop has been unable to machine these 
discs successfully and we are now looking at outside sources.  We are 
attempting to have a disc in time so that when the new window is installed 
the disc can be installed at the same time.

The new TripleSpec cart is now in routine operation.

Echelle is operational.  We have confirmed that the flip mirror is a 
source of variations in flats.  A number of fixes (both software and 
hardware) are in work and are being tested.  The problem is compounded by 
minimal documentation and the very old C+ software used for the motor 
control computer.

NIC-FPS is operating in shared-risk observing mode due to occasional image 
corruption issues.  A report of a glow in part of the image in K band 
appears to have been caused by a loose wiring harness in the FOV that has 
been corrected.


SPIcam status:

This was discussed above.


Echelle throughput:

Jon said that there is nothing new to report, other than that new data has 
been taken in recent weeks that needs to be reduced.  There followed a 
discussion that noted how user traffic on the echelle has increased over 
the past two years.  Jon asked if there might be any recent echelle users 
that can shed light on the instrumental throughput, before and after the 
grating replacement.


Upcoming BoG meeting:

Suzanne thanked the committee members for providing science highlight 
charts for the annual report.  Rene discussed the agenda for the upcoming 
Board meeting, which will be the usual with the addition of discussion on 
new partners.  We're intact through 2014 with existing partners, after 
which we'll need to replace Princeton and Chicago.  Rene added that we are 
in discussion with a handful of interested candidate institutions. 
Suzanne mentioned that there is a potential selling point for new 
institutions in the proposed fiber connection between the 3.5-m telescope 
and the APOGEE H-band spectrograph -- an ATI funding proposal is in the 
works for next year.  John Bally and Scott Anderson both said that there 
is enthusiasm at their institutions for this new capability.  We need the 
other committee members to poll their users, also.  Bruce brought up a 
recent meeting held at Sunspot where Mark Klaene learned of the likelihood 
that NSO will leave Sunspot by 2017.  This would have major impact on our 
water supply and local aluminizing capabilities.

  It was decided to send the link to the annual reports to the Users 
Committee members, so that they are prepared to answer questions brought 
to them before the Board meeting by their institutional ARC 
representatives.  Rene said that if there are other Board-related comments 
or questions to contact him.


Need references for Nov 2010-present:

Rene is updating the 3.5-m telescope publications list for the past two 
years.  Committee members are reminded to tell users to get their 
publications on the list; contact Rene with updates.


Other business:

There was no other business.



[open from previous meeting]:

===> ACTION:  Send new ARC partner suggestions to Suzanne and Rene (before next meeting).

 	Status:  Open, committee members need to keep sending suggestions.  We need the names of institutions (of any size) and a named contact.

===> ACTION:  Science slides for BOG meeting should be sent to Suzanne ASAP.

 	Status:  Completed, closed.

===> ACTION:  Users committee members should discuss APOGEE fiber feed to 3.5m proposal with their users and forward feedback, and ideally name(s) of interested users who would be willing to participate in an ATI proposal to Suzanne, Mark and Fred Hearty.

 	Status:  Open.


Next meeting:  The next Users Committee phonecon will be on Monday, 3 
December at 8:30 AM Pacific Time.  The agenda and other materials will be 
sent to the committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1297 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to